# Fractal transference principles for subsets of $\mathbb{N}^{d}$ of positive density

### v1

@@ 3.2 Universal upper bounds via singular values and Dirichlet series
$c_{0}\prod_{i=1}^{n}a_{i,j}^{-2}\ \leq\ L_{j}(\omega_{n})\ \leq\ C_{0}\prod_{i=1}^{n}a_{i,j}^{-2}.$ (16)
@@
This is a standard fact for one–dimensional continued fractions with a tail alphabet. A direct reference is [22, Section 6, Remark, p. 5023], which records the coincidence of Hausdorff, box, and packing dimensions in this setting.
Alternatively, one may combine [22, Corollary 5.9] with the general identity $\dim_{P}=\dim_{B}$ for conformal iterated function systems ([21, Theorem 3.1]) and the discussion in [22, Section 6]. ∎
@@
We emphasize that the coincidence in Lemma 3.12 is not automatic for general conformal iterated function systems. In fact, Mauldin–Urbański [22, Theorem 6.2] exhibited many examples where Hausdorff dimension is strictly smaller than box/packing dimension: for every $p\geq 2$ there exists $q\geq 1$ such that if $l\geq q$ and $I_{l}:=\{n^{p}:n\geq l\}$, then the corresponding continued–fraction digit–restricted set satisfies
@@ 4 Moran Fractal Constructions in Higher Dimensions
This is the standard mass distribution principle; see [7, Proposition 4.2]. ∎
@@ 5.2 Extreme seed sets and their dimension
In what follows we construct seed sets inside $E_{d}^{\mathrm{vec}}$. Later, when restricting digits to a subset $S\subset\mathcal{C}_{K}$, we may further intersect with $\mathcal{E}_{\mathbb{N}^{d}}$ (or with $\mathcal{E}_{S}$) in order to enforce divergence of digits. We do not impose additional coordinatewise restrictions beyond what is required. In [24], the fractal transference principle is formulated on $\mathcal{E}$ for $d=1$; for consistency with that framework, we will also work on $\mathcal{E}_{\mathbb{N}^{d}}$ in the final arguments.
Given the set $S^{\ast}$ obtained in Lemma 5.2, we now construct Moran-type subsets of $E_{d}^{\mathrm{vec}}$ whose digit vectors lie in $S^{\ast}$ and grow along a prescribed scale. These sets will serve as extreme seed sets, achieving the maximal Hausdorff dimension permitted by the growth properties of $S$.
@@
$R_{t,L}(K):=\Bigl\{\mathbf{x}\in E_{d}^{\mathrm{vec}}:\mathbf{a}_{n}(\mathbf{x})\in K\cap\bigl\{\mathbf{v}\in\mathbb{N}^{d}:\ t^{n}\leq\|\mathbf{v}\|_{\infty}<Lt^{n}\bigr\}\ \text{for all }n\geq 1\Bigr\}.$
@@
In this balanced case, one may also redefine the seed set by restricting the ambient space to $\mathcal{E}_{\mathbb{N}^{d}}$, that is, by replacing $E_{d}^{\mathrm{vec}}$ with $\mathcal{E}_{\mathbb{N}^{d}}$ in the definition of $R_{t,L}(K)$, without affecting any of the above conclusions.
@@
and the seed set may equivalently be defined with ambient space $\mathcal{E}_{\mathbb{N}^{d}}$ in place of $E_{d}^{\mathrm{vec}}$, without affecting any of the conclusions.
@@
We claim that $F\subset R_{t,L}(S^{\ast})$. Fix $\mathbf{x}\in F$ and choose for each $n$ a word $\omega\in\Sigma_{n}^{\ast}$ such that $\mathbf{x}\in\overline{I_{\omega}}$. If some coordinate $x_{j}\in\mathbb{Q}$, then $x_{j}$ would have to be an endpoint of a one–dimensional cylinder of arbitrarily large order. However, cylinder endpoints are convergents $p_{n}/q_{n}$, and the corresponding denominators satisfy $q_{n}\to\infty$ (as $q_{n+1}=a_{n+1}q_{n}+q_{n-1}\geq q_{n}+q_{n-1}$). Hence a fixed rational number can be an endpoint for only finitely many orders, a contradiction. Thus $\mathbf{x}\in((0,1)\setminus\mathbb{Q})^{d}$.
For such $\mathbf{x}$, membership in $\overline{I_{\omega}}$ implies $\mathbf{x}\in I_{\omega}$, and the length-$n$ word $\omega\in\Sigma_{n}^{\ast}$ with $\mathbf{x}\in I_{\omega}$ is unique. Therefore these words are consistent in $n$ and define an infinite admissible digit sequence in $S^{\ast}$, so $\mathbf{x}\in R_{t,L}(S^{\ast})$.
@@ 5.3.2 Growth conditions
Let $(\varepsilon_{k})_{k\geq 1}$ be a decreasing sequence of positive numbers with $\varepsilon_{k}\downarrow 0$. We record the quantitative constraints we will impose (inductively) on $M_{k}$ (and later on $W_{k}$). These conditions are chosen to control how much the insertion can distort continued-fraction cylinders; compare [24, Section 3.4].
@@ 5.3.3 Almost Lipschitz property with explicit estimates
The hypotheses (35)–(36) are the natural higher–dimensional counterparts of the one–dimensional assumptions in [24, Proposition 3.5]. The proof follows the same strategy, with the only changes coming from the geometry of dd–dimensional cylinders (diameter and separation estimates).
@@
For $\mathbf{x}_{1},\mathbf{x}_{2}\in R_{t,L}(B,A,(M_{k}),(W_{k}))$, set $\mathbf{y}_{i}=f(\mathbf{x}_{i})\in R_{t,L}(B)$ and let $a_{n}(\mathbf{y}_{i})$ denote their $n$-th digits. Define
$n:=s(\mathbf{y}_{1},\mathbf{y}_{2}):=\min\{m\geq 0:a_{m+1}(\mathbf{y}_{1})\neq a_{m+1}(\mathbf{y}_{2})\}.$
We distinguish whether the first discrepancy occurs before or after the first insertion scale $M_{k}$.
Case 1: $n<M_{k}$. Here the difference occurs before any insertions. Both points lie among finitely many possible pairs of cylinders, and by Proposition 2.5 and Lemma 2.6 there exists $\delta_{0}>0$ such that $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\delta_{0}$. Since $\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq 1$, choosing $C_{\gamma}^{(1)}:=\delta_{0}^{-\gamma}$ gives
@@
$\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})\quad\text{for all }1\leq i\leq n+\sum_{j=1}^{q}|W_{j}|.$
Thus the first discrepancy for $(\mathbf{x}_{1},\mathbf{x}_{2})$ occurs at
$n^{\prime}:=n+\sum_{j=1}^{q}|W_{j}|+1,$
with $\mathbf{a}_{n^{\prime}}(\mathbf{x}_{1})=\mathbf{a}_{n+1}(\mathbf{y}_{1})\neq\mathbf{a}_{n+1}(\mathbf{y}_{2})=\mathbf{a}_{n^{\prime}}(\mathbf{x}_{2})$.
Lemma 2.6 gives
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{1}\,\operatorname{diam}(I(\mathbf{a}_{1}(\mathbf{x}_{1}),\dots,\mathbf{a}_{n^{\prime}}(\mathbf{x}_{1})))\cdot\max_{j\in\mathcal{D}}\frac{|a_{n+1}^{(j)}(\mathbf{y}_{2})-a_{n+1}^{(j)}(\mathbf{y}_{1})|}{a_{n+1}^{(j)}(\mathbf{y}_{1})a_{n+1}^{(j)}(\mathbf{y}_{2})},$
where $\mathcal{D}=\{1\leq j\leq d:a_{n+1}^{(j)}(\mathbf{x}_{1})\neq a_{n+1}^{(j)}(\mathbf{x}_{2})\}$. Since $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$ for distinct $\mathbf{v},\mathbf{w}\in B$ and $a_{n+1}^{(j)}(\mathbf{y}_{i})\leq Lt^{n+1}$,
$\max_{j\in\mathcal{D}}\frac{|a_{n+1}^{(j)}(\mathbf{y}_{2})-a_{n+1}^{(j)}(\mathbf{y}_{1})|}{a_{n+1}^{(j)}(\mathbf{y}_{1})a_{n+1}^{(j)}(\mathbf{y}_{2})}\geq(Lt^{n+1})^{-2}.$
@@ 6 Proofs of the main theorems
$\frac{\#(A_{2}\cap Q_{L_{j}})}{\#(A_{1}\cap Q_{L_{j}})}\longrightarrow\delta\qquad(j\to\infty).$ (43)
@@
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$ (44)
@@
$\frac{\#((\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}.$ (48)
@@
$\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(A_{2}\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}-\frac{\#(A_{1}\cap Q_{r_{k,\min}})}{\#(A_{1}\cap Q_{R_{k}})}>\delta-\frac{1}{k}-\frac{1}{k}$ (50)
@@
$\#I_{k}=\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr)^{d}-\bigl(Lt^{M_{k-1}}\bigr)^{d}\leq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{d}.$ (53)
@@
$\prod_{w\in I_{k}}\|w\|_{\infty}^{-2}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2\#I_{k}}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2^{d+1}M_{k}^{1/2}}$ (55)
@@
$\overline{d}\!\left(\bigcup_{k\geq 1}W_{k}\,\Bigm|\,S\right)=\overline{d}(A\setminus S^{\ast}\mid S)=\overline{d}(A\mid S).$ (61)
@@
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$ (66)
@@
$\frac{|(\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}}|}{|Q_{R_{k}}|}\geq\frac{|W_{k}|}{|Q_{R_{k}}|}.$ (70)
@@
$\frac{|W_{k}|}{|Q_{R_{k}}|}\geq\frac{|A^{(k)}\cap Q_{R_{k}}|}{|Q_{R_{k}}|}-\frac{|Q_{r_{k,\min}}|}{|Q_{R_{k}}|}>\delta-\frac{1}{k}-\frac{1}{k}.$ (72)
@@
$\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)=\delta\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S},$ (73)
@@
$\mathcal{F}_{S}:=R_{t,L}\bigl(B,\mathbb{N}^{d},(M_{k}),(W_{k})\bigr).$ (78)
@@
$W\subset D(\mathbf{x})\cap S\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S}.$ (79)
@@
$\overline{d}(S)=\overline{d}(W)\leq\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)\leq\overline{d}(S),$ (80)
@@
$\dim_{H}(\mathcal{F}_{S})\geq\dim_{H}(\mathcal{R})=\frac{d}{2}.$ (82)
@@
$\frac{|S_{0}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}\longrightarrow\beta.$ (83)
Replacing $S$ by $S_{0}$ (and relabelling), we may assume $\overline{d}(S)=0$ and (83) holds with $\beta=d^{\ast}(S)$.
Passing to a subsequence, we may assume the boxes $B_{k}:=Q_{N_{k}}(\mathbf{v}_{k})$ are pairwise disjoint. Define $W_{k}:=S\cap B_{k}$ and $m_{k}:=|W_{k}|$. Then (83) becomes
$\frac{|W_{k}|}{|Q_{N_{k}}|}\longrightarrow d^{\ast}(S)=\beta.$ (84)
@@
For each $k$, the construction (86) inserts the block $W_{k}$ (in a fixed order) at the digit positions $\{M_{k},\dots,M_{k}+m_{k}-1\}$, independently of $\mathbf{x}\in F_{S}$. Thus $W_{k}\subset D_{*}\cap S$ for every $k$, and therefore
@@
Taking $\limsup$ and using (84) gives $d^{\ast}(D_{*}\cap S)\geq d^{\ast}(S)$, while $d^{\ast}(D_{*}\cap S)\leq d^{\ast}(S)$ is immediate from $D_{*}\cap S\subset S$. Hence

### v2

@@ 3.2 Universal upper bounds via singular values and Dirichlet series
$c_{0}\prod_{i=1}^{n}(a_{i,j}+1)^{-2}\ \leq\ L_{j}(\omega_{n})\ \leq\ C_{0}\prod_{i=1}^{n}a_{i,j}^{-2}.$ (9)
@@
This is a standard fact for one–dimensional continued fractions with a tail alphabet. A direct reference is [31, Section 6, Remark, p. 5023], which records the coincidence of Hausdorff, box, and packing dimensions in this setting.
Alternatively, one may combine [31, Corollary 5.9] with the general identity $\dim_{P}=\dim_{B}$ for conformal iterated function systems ([30, Theorem 3.1]) and the discussion in [31, Section 6]. ∎
@@
We emphasize that the coincidence in Lemma 3.12 is not automatic for general conformal iterated function systems. In fact, Mauldin–Urbański [31, Theorem 6.2] exhibited many examples where Hausdorff dimension is strictly smaller than box/packing dimension: for every $p\geq 2$ there exists $q\geq 1$ such that if $l\geq q$ and $I_{l}:=\{n^{p}:n\geq l\}$, then the corresponding continued–fraction digit–restricted set satisfies
@@ 4 Moran Fractal Constructions in Higher Dimensions
This is the standard mass distribution principle; see [10, Proposition 4.2]. ∎
@@ 5.2 Extreme seed sets and their dimension
In what follows we construct seed sets inside $E_{\mathbb{N}^{d}}^{\mathrm{vec}}$. Later, when restricting digits to a subset $S\subset\mathcal{C}_{K}$, we may further intersect with $\mathcal{E}_{\mathbb{N}^{d}}$ (or with $\mathcal{E}_{S}$) in order to enforce divergence of digits. We do not impose additional coordinatewise restrictions beyond what is required. In [34], the fractal transference principle is formulated on $\mathcal{E}$ for $d=1$; for consistency with that framework, we will also work on $\mathcal{E}_{\mathbb{N}^{d}}$ in the final arguments.
Given the set $S^{\ast}$ obtained in Lemma 5.2, we now construct Moran-type subsets of $E_{\mathbb{N}^{d}}^{\mathrm{vec}}$ whose digit vectors lie in $S^{\ast}$ and grow along a prescribed scale. These sets will serve as extreme seed sets, achieving the maximal Hausdorff dimension permitted by the growth properties of $S$.
@@
$R_{t,L}(K):=\Bigl\{\mathbf{x}\in E_{\mathbb{N}^{d}}^{\mathrm{vec}}:\mathbf{a}_{n}(\mathbf{x})\in K\cap\bigl\{\mathbf{v}\in\mathbb{N}^{d}:\ t^{n}\leq\|\mathbf{v}\|_{\infty}<Lt^{n}\bigr\}\ \text{for all }n\geq 1\Bigr\}.$
@@
In this balanced case, one may also redefine the seed set by restricting the ambient space to $\mathcal{E}_{\mathbb{N}^{d}}$, that is, by replacing $E_{\mathbb{N}^{d}}^{\mathrm{vec}}$ with $\mathcal{E}_{\mathbb{N}^{d}}$ in the definition of $R_{t,L}(K)$, without affecting any of the above conclusions.
@@
and the seed set may equivalently be defined with ambient space $\mathcal{E}_{\mathbb{N}^{d}}$ in place of $E_{\mathbb{N}^{d}}^{\mathrm{vec}}$, without affecting any of the conclusions.
@@
We show $F\subset R_{t,L}(S^{\ast})$ by proving that every $\mathbf{x}\in F$ has irrational coordinates; in particular, $\mathbf{x}\notin\overline{I_{\omega}}\setminus I_{\omega}$ for all $\omega$. Therefore $\mathbf{x}\in R_{t,L}(S^{\ast})$.
Fix $\mathbf{x}\in F$. For each $n\geq 1$ there exists $\omega^{(n)}\in\Sigma_{n}^{\ast}$ such that $\mathbf{x}\in\overline{I_{\omega^{(n)}}}$. Suppose $x_{j}\in\mathbb{Q}$ for some $j$. Then $x_{j}$ has only finite continued–fraction expansions, hence a maximal length $L$.
On the other hand, $x_{j}\in\overline{I^{(j)}_{\omega^{(n)}}}$ for all $n$ implies that for each $n$, $x_{j}$ shares a continued–fraction prefix of length at least n1n-1 with $\omega^{(n)}$. Thus $x_{j}$ admits expansions of arbitrarily large length, a contradiction. Hence $x_{j}\notin\mathbb{Q}$ for every $j$. Consequently $\mathbf{x}$ has a unique infinite expansion compatible with $S^{\ast}$, and therefore $\mathbf{x}\in R_{t,L}(S^{\ast})$.
@@ 5.3.2 Growth conditions
Let $(\varepsilon_{k})_{k\geq 1}$ be a decreasing sequence of positive numbers with $\varepsilon_{k}\downarrow 0$. We record the quantitative constraints we will impose (inductively) on $M_{k}$ (and later on $W_{k}$). These conditions are chosen to control how much the insertion can distort continued-fraction cylinders; compare [34, Section 3.4].
@@ 5.3.3 Almost Lipschitz property with explicit estimates
The hypotheses (28)–(29) are the natural higher–dimensional counterparts of the one–dimensional assumptions in [34, Proposition 3.5]. The proof follows the same strategy, with the only changes coming from the geometry of dd–dimensional cylinders (diameter and separation estimates).
@@
For $\mathbf{x}_{1},\mathbf{x}_{2}\in R_{t,L}(B,A,(M_{k}),(W_{k}))$, set $\mathbf{y}_{i}:=f(\mathbf{x}_{i})\in R_{t,L}(B)$ and define
$n:=\min\{m\geq 0:\ \mathbf{a}_{m+1}(\mathbf{y}_{1})\neq\mathbf{a}_{m+1}(\mathbf{y}_{2})\}.$
Since $B$ has no nearest–neighbour pairs, equivalently
$n=\min\{m\geq 0:\ \|\mathbf{a}_{m+1}(\mathbf{y}_{1})-\mathbf{a}_{m+1}(\mathbf{y}_{2})\|_{\infty}\geq 2\}.$
Fix $k\in\mathbb{N}$ and distinguish whether $n<M_{k}$ or $n\geq M_{k}$.
Case 1: $n<M_{k}$. Then the discrepancy occurs before any insertions, so $(\mathbf{x}_{1},\mathbf{x}_{2})$ ranges over finitely many cylinder pairs. By Proposition 2.5 and Lemma 2.6 there exists $\delta_{0}=\delta_{0}(k)>0$ such that $\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq\delta_{0}$. Since $\|\mathbf{y}_{1}-\mathbf{y}_{2}\|_{\infty}\leq 1$, taking $C_{\gamma}^{(1)}:=\delta_{0}^{-\gamma}$ yields
@@
$\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})\qquad\text{for all }1\leq i\leq n+\sum_{j=1}^{q}|W_{j}|.$
Hence the first discrepancy for $(\mathbf{x}_{1},\mathbf{x}_{2})$ occurs at
$n^{\prime}:=n+\sum_{j=1}^{q}|W_{j}|+1.$
Moreover, insertions only shift indices, so $\mathbf{a}_{n^{\prime}}(\mathbf{x}_{i})=\mathbf{a}_{n+1}(\mathbf{y}_{i})$ for $i=1,2$. Since distinct vectors in $B$ satisfy $\|\mathbf{v}-\mathbf{w}\|_{\infty}\geq 2$, we have
$\mathbf{a}_{i}(\mathbf{x}_{1})=\mathbf{a}_{i}(\mathbf{x}_{2})\ \ (1\leq i\leq n^{\prime}-1),\qquad\|\mathbf{a}_{n^{\prime}}(\mathbf{x}_{1})-\mathbf{a}_{n^{\prime}}(\mathbf{x}_{2})\|_{\infty}=\|\mathbf{a}_{n+1}(\mathbf{y}_{1})-\mathbf{a}_{n+1}(\mathbf{y}_{2})\|_{\infty}\geq 2.$
Apply Lemma 2.6,
$\|\mathbf{x}_{1}-\mathbf{x}_{2}\|_{\infty}\geq C_{1}\max_{j\in\mathcal{D}}\left(|I(a_{1}^{(j)}(\mathbf{x}_{1}),\ldots,a_{n^{\prime}-1}^{(j)}(\mathbf{x}_{1}))|\cdot\frac{|a_{n^{\prime}}^{(j)}(\mathbf{x}_{1})-a_{n^{\prime}}^{(j)}(\mathbf{x}_{2})|}{a_{n^{\prime}}^{(j)}(\mathbf{x}_{1})\,a_{n^{\prime}}^{(j)}(\mathbf{x}_{2})}\right),$
where $\mathcal{D}:=\{1\leq j\leq d:\ |a_{n^{\prime}}^{(j)}(\mathbf{x}_{1})-a_{n^{\prime}}^{(j)}(\mathbf{x}_{2})|\geq 2\}$. Since $a_{n^{\prime}}(\mathbf{x}_{i})=a_{n+1}(\mathbf{y}_{i})\in B$ and $\mathbf{y}_{i}\in R_{t,L}(B)$, we have $a_{n^{\prime}}^{(j)}(\mathbf{x}_{i})\leq Lt^{\,n+1}$ for all $j$ and $i=1,2$, and hence
$\max_{j\in\mathcal{D}}\frac{|a_{n^{\prime}}^{(j)}(\mathbf{x}_{1})-a_{n^{\prime}}^{(j)}(\mathbf{x}_{2})|}{a_{n^{\prime}}^{(j)}(\mathbf{x}_{1})\,a_{n^{\prime}}^{(j)}(\mathbf{x}_{2})}=\max_{j\in\mathcal{D}}\frac{|a_{n+1}^{(j)}(\mathbf{y}_{1})-a_{n+1}^{(j)}(\mathbf{y}_{2})|}{a_{n+1}^{(j)}(\mathbf{y}_{1})\,a_{n+1}^{(j)}(\mathbf{y}_{2})}\geq(Lt^{\,n+1})^{-2}.$
@@ 6 Proofs of the main theorems
$\frac{\#(A_{2}\cap Q_{L_{j}})}{\#(A_{1}\cap Q_{L_{j}})}\longrightarrow\delta\qquad(j\to\infty).$
@@
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$
@@
$\frac{\#((\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}.$
@@
$\frac{\#(W_{k})}{\#(A_{1}\cap Q_{R_{k}})}\geq\frac{\#(A_{2}\cap Q_{R_{k}})}{\#(A_{1}\cap Q_{R_{k}})}-\frac{\#(A_{1}\cap Q_{r_{k,\min}})}{\#(A_{1}\cap Q_{R_{k}})}>\delta-\frac{1}{k}-\frac{1}{k}$
@@
$I_{k}=\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}\bigr)^{d}-\bigl(Lt^{M_{k-1}}\bigr)^{d}\leq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{d}.$
@@
$\prod_{w\in I_{k}}\|w\|_{\infty}^{-2}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2\#I_{k}}\geq\bigl(Lt^{M_{k-1}}+M_{k}^{1/(2d)}+1\bigr)^{-2^{d+1}M_{k}^{1/2}}$
@@
$\overline{d}\!\left(\bigcup_{k\geq 1}W_{k}\,\Bigm|\,S\right)=\overline{d}(A\setminus S^{\ast}\mid S)=\overline{d}(A\mid S).$
@@
$r_{k,\min}:=\bigl\lfloor Lt^{M_{k-1}}\bigr\rfloor\in\mathbb{N}.$
@@
$\frac{|(\bigcup_{m\geq 1}W_{m})\cap Q_{R_{k}}|}{|Q_{R_{k}}|}\geq\frac{|W_{k}|}{|Q_{R_{k}}|}.$
@@
$\frac{|W_{k}|}{|Q_{R_{k}}|}\geq\frac{|A^{(k)}\cap Q_{R_{k}}|}{|Q_{R_{k}}|}-\frac{|Q_{r_{k,\min}}|}{|Q_{R_{k}}|}>\delta-\frac{1}{k}-\frac{1}{k}.$
@@
$\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)=\delta\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S},$
@@
$\mathcal{F}_{S}:=R_{t,L}\bigl(B,\mathbb{N}^{d},(M_{k}),(W_{k})\bigr).$
@@
$W\subset D(\mathbf{x})\cap S\qquad\text{for every }\mathbf{x}\in\mathcal{F}_{S}.$
@@
$\overline{d}(S)=\overline{d}(W)\leq\overline{d}\bigl(D(\mathbf{x})\cap S\bigr)\leq\overline{d}(S),$
@@
$\dim_{H}(\mathcal{F}_{S})\geq\dim_{H}(\mathcal{R})=\frac{d}{2}.$
@@
$\frac{|S_{0}\cap Q_{N_{k}}(\mathbf{v}_{k})|}{|Q_{N_{k}}|}\longrightarrow\beta.$
Replacing $S$ by $S_{0}$ (and relabelling), we may assume $\overline{d}(S)=0$ and (6.3.2) holds with $\beta=d^{\ast}(S)$.
Passing to a subsequence, we may assume the boxes $B_{k}:=Q_{N_{k}}(\mathbf{v}_{k})$ are pairwise disjoint. Define $W_{k}:=S\cap B_{k}$ and $m_{k}:=|W_{k}|$. Then (6.3.2) becomes
$\frac{|W_{k}|}{|Q_{N_{k}}|}\longrightarrow d^{\ast}(S)=\beta.$
@@
For each $k$, the construction (6.3.2) inserts the block $W_{k}$ (in a fixed order) at the digit positions $\{M_{k},\dots,M_{k}+m_{k}-1\}$, independently of $\mathbf{x}\in F_{S}$. Thus $W_{k}\subset D_{*}\cap S$ for every $k$, and therefore
@@
Taking $\limsup$ and using (6.3.2) gives $d^{\ast}(D_{*}\cap S)\geq d^{\ast}(S)$, while $d^{\ast}(D_{*}\cap S)\leq d^{\ast}(S)$ is immediate from $D_{*}\cap S\subset S$. Hence